OpenStudy (anonymous):

The area of a hall is 806 sq. ft. Its length is 5 ft greater than its width. Find the dimensions of the hall.

OpenStudy (anonymous):

let the width be = x ft then length will be = (x + 5) ft Area of rectangle = length * breadth/width so 806 = (x + 5)(x) 806 = x^2 + 5x so we get the quadratic equation x^2 + 5x - 806 = 0 this will factor out to (x + 31)(x - 26) so x = - 31 or x = 26 since lengths are never negative, so x = 26 is the acceptable value. so width = x = 26 ft and width = x + 5 = 26 + 5 = 31 ft

OpenStudy (anonymous):

@Happysoul wait l thought area of a rect is l*w

OpenStudy (anonymous):

yes you are right, area of rectangle is the product of its length (l) and its width (w) i.e. l*w or lw I have solved it on that basis only.......☺

OpenStudy (anonymous):

@HappySoul can you help me with this one The area of a hall is 806 sq. ft. Its length is 5 ft greater than its width. Find the dimensions of the hall.

OpenStudy (anonymous):

sorry not that one

OpenStudy (anonymous):

I have given the solution above......This is the same question.....

OpenStudy (anonymous):

It took a motorboat 1 hour longer to travel 84 miles against the current than travel the same distance with the current. The speed of the current is 1 mph. What is the speed of the motorboat in still water?

OpenStudy (anonymous):

****In such questions, you have to add the speed of the current to the speed of boat while going downstream and subtract while going upstream**** let the speed of motorboat in still water = x mph then when it goes down the stream, its speed will be = (x+1) mph and speed while going up the stream will be = (x - 1) mph distance in both cases is 84 miles let the time while going downstream be = T1 and time while going upstream be = T2 Time = distance/speed so 84 84 T1= ----- and T2 = ------ x+1 x-1 according to the question the boat takes 1 more hour while going upstream (against the current) so T2 = T1 + 1 84 84 ----- = ----- + 1 x - 1 x +1 84 84 + (x + 1) ----- = ------------- x - 1 (x + 1) 84 85 + x ----- = -------- x - 1 (x + 1) cross multiplying 84(x + 1) = (85 + x)(x - 1) simplify this to get a quadratic equation and solve it to get your answer